Problem 2. We have 4 urns and for each k = 1, 2, 3, 4, um k contains k red and 10-k green balls. We choose an urn randomly and then draw a ball from it at random. What is the probabláty that we draw a red ball?
Answer #1

probability that we draw red ball =P(1st urn and red ball+2nd urn and red ball+3rd urn and red ball+4th urn and red ball) =(1/4)*(1/10)+(1/4)*(2/10)+(1/4)*(3/10)+(1/4)*(4/10)=1/4 =0.25